Please solve the system of equations by the method of your choice:
2x-4y=6 and 4x+y=21
Are these equations independent, inconsistent, or dependent?
Thank you,
Shelley E.

Answer by stanbon(75887)

Please solve the system of equations by the method of your choice:
2x-4y=6
4x+y=21
Are these equations independent, inconsistent, or dependent?
----------------------
Multiply the 1st equation by 2:
4x-8y = 12
4x+y = 21
----------------
Subtract the 1st from the 2nd and solve for "y":
9y = 9
y = 1
---
Substitute into 2x-4y = 6 to solve for "x":
2x-4*1 = 6
2x = 10
x = 5
-------------
Cheers,
Stan H.
===========